b1az3:

Solve the following system. y = -2x 2 + 3x - 1 and x + y = -1. The solutions are and

imqwerty:

Where are you stuck?

b1az3:

yes

imqwerty:

no

jhonyy9:

y = -2x^2 +3x -1 x+y = -1 what method you wan - need using to solve it ? substitution ,elimination,...?

b1az3:

it just askes this Solve the following system. y = -2x 2 + 3x - 1 and x + y = -1. The solutions are and

jhonyy9:

so the most easy way to use it the substitution method bc. there given a value of y what you can substitute it inside the second equation in place of y and solve it for x

imqwerty:

^

jhonyy9:

@AZ any idea ? ty

jhonyy9:

@b1az3 do you understand the way ?

b1az3:

no

jhonyy9:

y = -2x^2 +3x -1 this expression of y you need substitute it in place of y inside second equation

jhonyy9:

so what will get ?

jhonyy9:

@imqwerty do you have any idea ?

jhonyy9:

x +y = -1 x +(-2x^2 +3x -1) = -1 solve this for x

imqwerty:

^